# Tunneling in double-well potentials within stochastic quantization: Application to ammonia inversion

### v1

@@ VI Discussion
In this work, we have employed Nelson’s stochastic quantization (SQ) to investigate tunneling-time statistics for bound states in symmetric double-well potentials. By combining the stochastic-trajectory framework with the theory of first-passage times, we obtained closed expressions for both the mean tunneling time, $\bar{\tau}$ and the asymptotic (tail) behavior of the probability distribution $p(\tau)$. For the double square-well potential, an analytic formula was derived for $\bar{\tau}$ and shown to be in excellent agreement with large-scale numerical simulations of the underlying diffusion process.
@@
Figure 10: Ratio $\tau_{\rm QM}/\bar{\tau}$ as a function of $x_{f}$ for the Rosen-Morse potential, Eq. (84), with $V_{0}=\{39.5\text{ meV},\,98.0\text{ meV},\,286.5\text{ meV}\}$ corresponding to points A, B and C in Fig. 9. The $x_{f}$ parameter, which delimits the tunneling region, varies from the classical turning point on the barrier ($b$) to the minimum potential point ($x_{0}$). If $V_{0}$ is high, then the ratio gets close to the value $\pi/2$ despite the choice of the parameter $x_{f}$.
@@
In Fig. 10 we plot $\bar{\tau}(x_{f})$ as a function of the initial tunneling point, $x_{f}$, for the RMP potential, using parameters corresponding to points A, B, and C in Fig. 9. For the three curves shown, $x_{f}$ is varied from the minimum point of the left well ($x_{0}$) to the right classical turning point (b-b). As expected, for a low barrier (black circles), the ratio $\tau_{\mathrm{QM}}/\bar{\tau}$ changes more significantly with $x_{f}$. As the barrier height increases—see, for instance, the upper curve in Fig. 10—this ratio becomes insensitive to the precise definition of the tunneling stopping rule, thereby confirming the universal character of the factor $\pi/2$ in this limit. This result provides further evidence that the classical turning points on both sides of the barrier can be safely used as reference positions in defining the quantum tunneling process, as argued above.
@@
Physically, the proportionality between $\tau_{\mathrm{QM}}$ and $\bar{\tau}$ indicates that the diffusive motion intrinsic to stochastic quantization reproduces the same characteristic timescale that, in quantum mechanics, governs coherent oscillations between the two localized states of the double-well. However, the mean tunneling time in SQ is computed solely within the ground state and requires no prior knowledge of the energy splitting—a quantity that is essential for describing tunneling in standard quantum mechanics. The fact that $\tau_{\mathrm{QM}}$ exceeds $\bar{\tau}$ may be understood as reflecting the distinction between two complementary descriptions of the tunneling process. In the stochastic picture, $\bar{\tau}$ represents the mean first-passage time for an individual trajectory to cross the barrier from one well to the other, whereas in quantum mechanics $\tau_{\mathrm{QM}}$ relates to the collective oscillation period of the entire probability distribution. The latter entails a coherent redistribution of probability density between the wells—a process inherently slower than the stochastic transition of single trajectories. Within the high-barrier, two-level approximation, this interpretation provides a consistent physical link between single-trajectory diffusion dynamics and the collective evolution of quantum probability. (Whether a physical explanation exists for the emergence of $\pi/2$ as the universal ratio remains unclear at present.)
@@
As discussed above, the SQ formulation offers several conceptual advantages for studying dynamical properties of quantum systems. Because SQ is expressed directly in real space and time, it allows one to define operational quantities—such as traversal-time distributions—that are difficult or ambiguous within the standard Hilbert-space framework, where time appears only as an external parameter. The emergence of exponentially distributed tunneling times, derived analytically in general and confirmed numerically for both the square and Rosen–Morse double-wells, indicates that stochastic trajectories effectively sample barrier-crossing events as a Poisson-like process characterized by a single mean scale $\bar{\tau}$. This provides a concrete dynamical interpretation of tunneling as a rare-event diffusion process underlying the coherent oscillations described by quantum mechanics.
@@
The successful application of the SQ formalism to the inversion dynamics of the ammonia molecule further illustrates its predictive power. Using the double Rosen–Morse potential and the formula for the stochastic-mechanical tunneling frequency $\nu_{\rm SQ}$ in terms of $\bar{\tau}$, see Eq.(83), we obtained an inversion frequency for ammonia of about $24~\mathrm{GHz}$, in remarkable agreement with the experimental value. This quantitative concordance demonstrates that the SQ framework captures not only the qualitative behavior of tunneling but also its measurable spectroscopic signature. It also reinforces the view that the stochastic representation is not merely an interpretational alternative but a mathematically equivalent and physically transparent formulation of quantum mechanics, offering practical computational advantages in certain contexts.
@@ VII Conclusions
The present work demonstrates that Nelson’s stochastic quantization provides a coherent and quantitatively accurate framework for describing quantum tunneling as a diffusive process in real space and time. By establishing an explicit link between the mean stochastic tunneling time and the quantum‐mechanical oscillation period, we have shown that SQ can reproduce measurable quantum timescales while offering a more intuitive dynamical picture based on stochastic trajectories. The derived relation $\tau_{\mathrm{QM}}=(\pi/2)\,\bar{\tau}$, validated both analytically and numerically, reveals a universal correspondence between the stochastic and quantum descriptions in the high‐barrier limit.
The stochastic-quantization formalism has already been extended to several important cases beyond the one-dimensional systems studied here. These include dissipative quantum systems [Marra1987, Bacciagaluppi2012], multi-particle systems exhibiting quantum nonlocality [delaPena1969, Loffredo2007], and semiclassical dynamics [JonaLasinio1981]. A comprehensive modern treatment of stochastic mechanics, including relativistic extensions and applications on curved spacetimes, has been recently provided by Kuipers [Kuipers2023]. By providing direct access to first-passage-time distributions—quantities not easily obtained from the standard quantum formalism—stochastic quantization may offer complementary insights into tunneling processes in complex molecular and condensed-matter systems.
@@ Appendix A Unitary Equivalence to Quantum Mechanics
While the hydrodynamical equivalence discussed in Sec. II establishes the correspondence between Nelson’s stochastic quantization and quantum mechanics through the Madelung flow equations, we can gain deeper insight by examining the operator formulation of this equivalence. This alternative perspective reveals that the stochastic and quantum evolutions are actually related through a unitary transformation, which provides a more profound understanding of the mathematical structure underlying the correspondence between the two formalisms.
The key observation here is that the forward and backward stochastic processes can be combined into a single complex-valued drift, leading to what is called a bi-directional generator [Pavon2000]. To see how this works, let us first introduce a complex-valued quantum drift velocity defined by
@@
The central result that establishes the operator equivalence can be stated as follows [Pavon2000]. Let $H=-\frac{\hbar^{2}}{2m}\Delta+V(\mathbf{x})$ be the Hamiltonian operator and $\psi(\mathbf{x},t)$ be a solution of the Schrödinger equation
@@
We assume throughout that $\psi$ is never vanishing on the time interval considered and satisfies the finite-action condition $\int\!\!\int|\nabla\psi|^{2}\,dx\,dt<\infty$. Under these conditions, one can show [Pavon2000] that the stochastic evolution operator $\partial_{t}+L_{b}$ is unitarily equivalent to the quantum evolution operator $\partial_{t}+\frac{i}{\hbar}H$ via the conjugation
@@
It is worth noting that if $\psi$ has nodes (zeros), the equivalence still persists in a more general sense through regularization techniques. However, for most practical purposes, working with nodeless wavefunctions is sufficient. This is truly a remarkable result: it shows that the stochastic and quantum descriptions are mathematically identical up to a unitary transformation. The operator $M_{\psi}$ serves as the bridge between the two formulations, transforming quantum mechanical objects into their stochastic counterparts and vice versa.

### v2

@@ VI Discussion
In this work, we have employed Nelson’s stochastic mechanics (SM) to investigate tunneling-time statistics for bound states in symmetric double-well potentials. By combining the stochastic-trajectory framework with the theory of first-passage times, we obtained closed expressions for both the mean tunneling time, $\bar{\tau}$ and the asymptotic (tail) behavior of the probability distribution $p(\tau)$. For the double square-well potential, an analytic formula was derived for $\bar{\tau}$ and shown to be in excellent agreement with large-scale numerical simulations of the underlying diffusion process.
@@
Figure 10: Ratio $\tau_{\rm QM}/\bar{\tau}$ as a function of $x_{f}/d$ for the Rosen-Morse potential, Eq. (86), with $V_{0}=\{39.5\text{ meV},\,98.0\text{ meV},\,286.5\text{ meV}\}$ corresponding to points A, B and C in Fig. 9. The $x_{f}$ parameter, which delimits the tunneling region, varies from the classical turning point on the barrier ($b$) to the minimum potential point ($x_{0}$). If $V_{0}$ is high, then the ratio gets close to the value $\pi/2$ despite the choice of the parameter $x_{f}$.
@@
In Fig. 10 we plot the ratio $\tau_{\rm QM}/\bar{\tau}$ as a function of the initial tunneling point, $x_{f}$, for the RMP potential, using parameters corresponding to points A, B, and C in Fig. 9. For the three curves shown, $x_{f}$ is varied from the minimum point of the left well ($x_{0}$) to the right classical turning point (b-b). Since $\tau_{\mathrm{QM}}$ is fixed for each curve, one sees from Fig. 10 that $\bar{\tau}$ decreases as $x_{f}$ decreases, as expected, since the distance traveled within the tunneling region $(-x_{f},x_{f})$ becomes smaller. Also as expected, for a low barrier (black circles), the ratio $\tau_{\mathrm{QM}}/\bar{\tau}$ changes more significantly with $x_{f}$. As the barrier height increases—see, for instance, the upper curve in Fig. 10—this ratio becomes insensitive to the precise definition of the tunneling stopping rule, thereby confirming the universal character of the factor $\pi/2$ in this limit. This result provides further evidence that the classical turning points on both sides of the barrier can be safely used as reference positions in defining the quantum tunneling process, as argued above.
@@
Physically, the proportionality between $\tau_{\mathrm{QM}}$ and $\bar{\tau}$ indicates that the diffusive motion intrinsic to stochastic mechanics reproduces the same characteristic timescale that, in quantum mechanics, governs coherent oscillations between the two localized states of the double-well. However, the mean tunneling time in SM is computed solely within the ground state and requires no prior knowledge of the energy splitting—a quantity that is essential for describing tunneling in standard quantum mechanics. The fact that $\tau_{\mathrm{QM}}$ exceeds $\bar{\tau}$ may be understood as reflecting the distinction between two complementary descriptions of the tunneling process. In the stochastic picture, $\bar{\tau}$ represents the mean first-passage time for an individual trajectory to cross the barrier from one well to the other, whereas in quantum mechanics $\tau_{\mathrm{QM}}$ relates to the collective oscillation period of the entire probability distribution. The latter entails a coherent redistribution of probability density between the wells—a process inherently slower than the stochastic transition of single trajectories. Within the high-barrier, two-level approximation, this interpretation provides a consistent physical link between single-trajectory diffusion dynamics and the collective evolution of quantum probability. (Whether a physical explanation exists for the emergence of $\pi/2$ as the universal ratio remains unclear at present.)
@@
As discussed above, the SM formulation offers several conceptual advantages for studying dynamical properties of quantum systems. Because SM is expressed directly in real space and time, it allows one to define operational quantities—such as traversal-time distributions—that are difficult or ambiguous within the standard Hilbert-space framework, where time appears only as an external parameter. The emergence of exponentially distributed tunneling times, derived analytically in general and confirmed numerically for both the square and Rosen–Morse double-wells, indicates that stochastic trajectories effectively sample barrier-crossing events as a Poisson-like process characterized by a single mean scale $\bar{\tau}$. This provides a concrete dynamical interpretation of tunneling as a rare-event diffusion process underlying the coherent oscillations described by quantum mechanics.
@@
The successful application of the SM formalism to the inversion dynamics of the ammonia molecule further illustrates its predictive power. Using the double Rosen–Morse potential and the formula for the stochastic-mechanical tunneling frequency $\nu_{\rm SM}$ in terms of $\bar{\tau}$ (see Eq. (85)), we obtained an inversion frequency of approximately $24~\mathrm{GHz}$, in remarkable agreement with the experimental value. This quantitative concordance demonstrates that the SM framework captures not only the qualitative features of tunneling dynamics but also its measurable spectroscopic signature. In this sense, the stochastic representation is not merely an interpretational alternative but a dynamically consistent formulation of quantum mechanics at the level of Schrödinger evolution and single-time observables, while offering practical computational advantages in certain contexts.
The successful application of the SM formalism to the inversion dynamics of the ammonia molecule further illustrates its predictive power. Using the double Rosen–Morse potential and the formula for the stochastic-mechanical tunneling frequency $\nu_{\rm SM}$ in terms of $\bar{\tau}$, see Eq.(85), we obtained an inversion frequency for ammonia of about $24~\mathrm{GHz}$, in remarkable agreement with the experimental value. This quantitative concordance demonstrates that the SM framework captures not only the qualitative behavior of tunneling but also its measurable spectroscopic signature. In this sense, the stochastic representation is not merely an interpretational alternative but a dynamically consistent formulation of quantum mechanics at the level of wavefunction evolution and associated single-time statistics, while offering practical computational advantages in certain contexts.
@@ VII Conclusions
The present work demonstrates that Nelson’s stochastic mechanics provides a coherent and quantitatively accurate framework for describing quantum tunneling as a diffusive process in real space and time. By establishing an explicit link between the mean stochastic tunneling time and the quantum‐mechanical oscillation period, we have shown that SM can reproduce measurable quantum timescales while offering a more intuitive dynamical picture based on stochastic trajectories. The derived relation $\tau_{\mathrm{QM}}=(\pi/2)\,\bar{\tau}$, validated both analytically and numerically, reveals a universal correspondence between the stochastic and quantum descriptions in the high‐barrier limit.
The stochastic-mechanics formalism has already been extended to several important cases beyond the one-dimensional systems studied here. These include dissipative quantum systems [38, 2], multi-particle systems exhibiting quantum nonlocality [11, 35], and semiclassical dynamics [25]. A comprehensive modern treatment of stochastic mechanics, including relativistic extensions and applications on curved spacetimes, has been recently provided by Kuipers [29]. By providing direct access to first-passage-time distributions—quantities not easily obtained from the standard quantum formalism—stochastic mechanics may offer complementary insights into tunneling processes in complex molecular and condensed-matter systems.
@@ Appendix A Unitary Equivalence to Quantum Mechanics
While the hydrodynamic equivalence discussed in Sec. II establishes a correspondence between Nelson’s stochastic mechanics framework and quantum mechanics through the Madelung flow equations, further insight can be gained by examining the operator formulation of this equivalence. This alternative perspective reveals that the stochastic and quantum evolutions are actually related through a unitary transformation, which provides a more profound understanding of the mathematical structure underlying the correspondence between the two formalisms.
The key observation here is that the forward and backward stochastic processes can be combined into a single complex-valued drift, leading to what is called a bi-directional generator [52]. To see how this works, let us first introduce a complex-valued quantum drift velocity defined by
@@
The central result that establishes the operator equivalence can be stated as follows [52]. Let $H=-\frac{\hbar^{2}}{2m}\Delta+V(\mathbf{x})$ be the Hamiltonian operator and $\psi(\mathbf{x},t)$ be a solution of the Schrödinger equation
@@
We assume throughout that $\psi$ is never vanishing on the time interval considered and satisfies the finite-action condition $\int\!\!\int|\nabla\psi|^{2}\,dx\,dt<\infty$. Under these conditions, one can show [52] that the stochastic evolution operator $\partial_{t}+L_{b}$ is unitarily equivalent to the quantum evolution operator $\partial_{t}+\frac{i}{\hbar}H$ via the conjugation
@@
It is worth noting that when the wavefunction $\psi$ possesses nodes (zeros), additional care is required, since the osmotic velocity involves gradients of $\log|\psi|$, which become singular on the nodal set. In such cases, the configuration space may be decomposed into nodal domains, and the correspondence between stochastic mechanics and the Schrödinger formulation can be extended through appropriate regularization procedures. For most practical purposes, however—as in the examples considered in the main text—, working with nodeless wavefunctions suffices, in which case the mapping between the two formulations is technically straightforward. In this setting, the operator $M_{\psi}$ establishes a unitary correspondence between the Hilbert-space representation and its stochastic counterpart.